# Exponential Runge–Kutta methods of collocation type for parabolic equations with time-dependent delay

### v1

@@ Abstract
In this paper, we extend the exponential Runge–Kutta methods of collocation type (ERKC) proposed in (Appl Numer Math 53:323–339, 2005) to semilinear parabolic problems with time-dependent delay. We construct two classes of the ERKC methods and analyze their convergence properties. It is shown that methods with $s$ arbitrary nonconfluent collocation parameters achieve convergence of order $s$. Provided that the collocation parameters fulfill some additional conditions and the solutions of the problems exhibit sufficient temporal and spatial smoothness, we derive superconvergence results of order $s+1$ and $s+1+\beta$. Finally, some numerical experiments are presented to illustrate our theoretical results.
@@ 1 Introduction
[…]
@@
In the past, significant numerical research has been conducted for problems of the form (1.1) with constant delay; see, for example, [8, 15, 21, 25]. Compared to constant delay models, time-dependent delay models generally provide a more accurate description of the dynamic nature of real-world systems. Nevertheless, the presence of time-dependent delays complicates both the mathematical modeling and the numerical solution, and relatively little numerical analysis has been performed for these models. Despite these difficulties, some studies have been conducted in this area. For example, Tang and Zhang [20, 29] constructed fully discrete $\theta$-methods to solve semilinear reaction-diffusion equations and Sobolev equations with time-dependent delay. However, their approach requires the absence of discontinuities in the true solution, which is a significant limitation, as discontinuities commonly arise in delay differential equations. Until now, there have been limited numerical studies on temporal high-order schemes for solving (1.1) that can account for discontinuities due to the presence of a time-dependent delay.
@@
In recent years, exponential integrators have attracted considerable attention due to their effectiveness on stiff semilinear systems, such as (1.1) with $A$ having a large norm or being an unbounded operator, and the nonlinearity $g$ satisfying a local Lipschitz condition with a relatively small Lip-schitz constant in a strip along the exact solution. The stiff convergence of exponential integrators, such as exponential Runge–Kutta methods [10, 11, 17], exponential Rosenbrock methods [14] and exponential multistep methods [13], has been investigated. Exponential integrators have also been extended to various areas, such as low-regularity integrators [19], exponential splitting methods [7], and phase field models [16]. For a detailed overview of exponential integrators, we refer the reader to [12]. Some research on exponential integrators for (parabolic) differential equations with constant delay can be found in [2, 4, 26, 27, 28, 30, 31]. In this paper, we aim to extend the exponential Runge–Kutta methods of collocation type (ERKC) as presented in [11] to the larger class of problems (1.1). For this purpose, we construct continuous extensions of the discrete solutions obtained by the ERKC methods, and we derive the convergence results.
@@
The outline of the paper is as follows. In Section 2, we summarize the employed abstract framework. Further, we construct two classes of $s$-stage ERKC methods for (1.1) in Section 3. The error analysis for these methods is carried out in Section 4, and an error estimate of order $s$ is obtained. Provided that the collocation parameters satisfy some additional conditions and the problem possesses sufficient temporal and spatial regularity, we obtain superconvergence results of order $s+1$ and $s+1+\beta$ in Section 5. Finally, we present some numerical experiments in Section 6 to illustrate the theoretical results.
@@ 2 Analytical Framework
Let $A:D(A)\rightarrow X$ be sectorial. Without restriction of generality, we assume that the resolvent set $\rho(A)$ of $A$, satisfies $\mathbb{C}^{-}\subset\rho(A)$, where =$\mathbb{C}^{-}=\{\lambda\in\mathbb{C}:\mathrm{Re}\,\lambda<0\}$.
Under this assumption, the operator $-A$ is the infinitesimal generator of an analytic semigroup $\{\mathrm{e}^{-tA}\}_{t\geq 0}$, and the fractional powers of $A$ are well defined. We recall that $A$ satisfies the parabolic smoothing property (see [9])
$\|\mathrm{e}^{-tA}\|_{X\leftarrow X}+\big{\|}t^{\gamma}A^{\gamma}\mathrm{e}^{-tA}\big{\|}_{X\leftarrow X}\leq C,\quad\gamma,t\geq 0.$
For $10\leq\alpha<$ 1, we consider the Banach space $V=\{v\in X:A^{\alpha}v\in X\}$ with norm $\|v\|_{V}=\|A^{\alpha}v\|$. Our basic assumption on $g$ is the following.
@@
[…]
@@
    $X=C(\Omega)$, $D(A)=\{u:u\in W^{2,p}(\Omega)~{}\mbox{for all}~{}p\geq 1,~{}\Delta u\in C(\Omega),~{}u|_{\partial\Omega}=0\}$.
@@ 3 ERKC methods
An important feature of delay partial differential equations is that they possess solutions with discontinuities in the derivative at initial time and discontinuities in high order derivatives at subsequent points. The set of potential discontinuity points $J=\{\xi_{\mu}:\mu\in\mathbb{N},~{}\xi_{\mu}<T\}$ (which are called primary discontinuity points in the theory of delay differential equations [1]) are generated by the recursion
@@
For $\xi_{m}<T\leq\xi_{m+1}$ with some $m\in\mathbb{N}$, we have $J=\{\xi_{\mu}\}_{\mu=0}^{m}$. At these points, the solutions of (1.1), regardless of how regular the given functions $\phi$, $g$ and $\tau$ are, will generally exhibit a low degree of regularity. Considering such discontinuity points, we choose the mesh
@@
such that $J\subseteq I_{h}$. We denote $\mathcal{I}^{\mu}=[\xi_{\mu-1},\min\{\xi_{\mu},T\}]$ for $11\leq\mu\leq m+$ 1. Choosing nonconfluent collocation parameters $c_{i}\in[0,1]$ for $i=1,2\ldots,s$, we set
$I_{n+1}=[t_{n},t_{n+1}],\quad h_{n+1}=t_{n+1}-t_{n},\quad t_{n,i}=t_{n}+c_{i}h_{n+1}\quad\mbox{and}\quad h=\max_{1\leq n\leq N}h_{n}.$
@@
$u(t_{n}+h_{n+1})=\mathrm{e}^{-h_{n+1}A}u(t_{n})+\int_{0}^{h_{n+1}}\mathrm{e}^{-(h_{n+1}-\sigma)A}g(t_{n}+\sigma,u(t_{n}+\sigma),u^{\tau}(t_{n}+\sigma))~{}\mathrm{d}\sigma.$ (3.3)
@@
$G_{n}(\sigma)=\sum_{j=1}^{s}\ell_{i}(\sigma/h_{n+1})G_{n,i},\quad\mbox{with}~{}G_{n,i}=g(t_{n,i},U_{n,i},u^{\tau}(t_{n,i})),$
where $U_{n,i}$ is an approximation to $u(t_{n,i})$ which will be determined later, and $\ell_{i}$ is the Lagrange interpolation polynomial
$\ell_{i}(\rho)=\prod_{m=1,m\neq i}^{s}\frac{\rho-c_{m}}{c_{i}-c_{m}}.$
@@
$b_{i}(-h_{n+1}A)=\frac{1}{h_{n+1}}\int_{0}^{h_{n+1}}\mathrm{e}^{-(h_{n+1}-\sigma)A}\ell_{i}(\sigma/h_{n+1})~{}\mathrm{d}\sigma.$ (3.5)
@@
$a_{ij}(-h_{n+1}A)=\frac{1}{h_{n+1}}\int_{0}^{c_{i}h_{n+1}}\mathrm{e}^{-(c_{i}h_{n+1}-\sigma)A}\ell_{j}(\sigma/h_{n+1})~{}\mathrm{d}\sigma.$ (3.7)
The scheme (3.4)-(3.7) is not computable yet, since the delayed values $u^{\tau}(t_{n,i})$ in $G_{n}$ still have to be approximated. One feasible approach is to construct a continuous approximation by using piecewise interpolation polynomials based on the known values $U_{k}$ and $U_{k,i}$ for $k=0,1,\ldots,n-1$ and $i=1,2,\ldots,s$. For this, we introduce an interpolation $\Pi_{h}U(t)$, which is a polynomial of degree $l$ when restricted to each interval $I_{k+1}$. This polynomial is based on the nodes of the set $S_{k+1}=\{t_{k},t_{k+1}\}\cup\{t_{k,1},t_{k,2},\ldots,t_{k,s}\}$, satisfying
$\Pi_{h}U(t_{k})=U_{k},\quad\Pi_{h}U(t_{k+1})=U_{k+1},\quad\mbox{and}\quad\Pi_{h}U(t_{k,i})=U_{k,i}~{}~{}\mbox{for}~{}i=1,2,\ldots,s.$ (3.8)
Depending on the number of distinct interpolation points available, the degree $l$ is either $s-1$, $s$ or $s+1$. For $t\in[-\tau_{0},0]$, we denote $\Pi_{h}U(t)=\phi(t)$. Replacing the evaluations $u^{\tau}(t_{n,i})$ in $G_{n}$ by $(\Pi_{h}U)^{\tau}(t_{n,i})$, we arrive at the scheme
@@
$U_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}U_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm I}_{n,j},\quad i=1,2,\ldots,s,$
$G^{\rm I}_{n,i}$ $=g(t_{n,i},U_{n,i},(\Pi_{h}U)^{\tau}(t_{n,i})),\quad i=1,2,\ldots,s,$
@@
The scheme (3.9) is one option among several possible schemes for extending ERKC methods to (1.1). We shall introduce an alternative approach for constructing an approximation to the delayed solution $u^{\tau}(t_{n,i})$ in $G_{n}$. To distinguish it from the previous introduced numerical solution $U_{n+1}$ and $U_{n,i}$ in (3.9), we denote the approximate solutions of the new method by $V_{n+1}$ and $V_{n,i}$:
$V_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}V_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(-h_{n+1}A)g(t_{n,i},V_{n,i},V^{\tau}(t_{n,i})),$
$V_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}V_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)g(t_{n,j},V_{n,j},V^{\tau}(t_{n,j})),\quad i=1,2,\ldots,s,$
with $V(t)$ still to be defined. Instead of using piecewise interpolation polynomials, we construct a continuous solution $V(t)$ by setting $V(t)=\phi(t)$ for $t\in[-\tau_{0},0]$ and, for $t=t_{k-1}+\theta h_{k}$ with some $k\in\{1,2,\ldots,n\}$ and $\theta\in(0,1]$,
$V(t_{k}+\theta h_{k+1})=\mathrm{e}^{-\theta h_{k+1}A}V_{k}+h_{k+1}\sum_{i=1}^{s}b_{i}(\theta;-h_{k+1}A)g(t_{k,i},V_{k,i},V^{\tau}(t_{k,i})),$ (3.10)
@@
$b_{i}(\theta;-h_{k+1}A)=\frac{1}{h_{k+1}}\int_{0}^{\theta h_{k+1}}\mathrm{e}^{-(\theta h_{k+1}-\sigma)A}\ell_{i}(\sigma/h_{k+1})~{}\mathrm{d}\sigma.$ (3.11)
@@
$V_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}V_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(-h_{n+1}A)G^{\rm C}_{n,i},$ (3.12)
$V_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}V_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm C}_{n,j},\quad i=1,2,\ldots,s,$
$G^{\rm C}_{n,i}$ $=g(t_{n,i},V_{n,i},V^{\tau}(t_{n,i})),\quad i=1,2,\ldots,s.$
For $A=0$, this scheme reduces to the collocation method, in which the delayed solution is approximated by the collocation polynomials. For this reason, we refer to it as the ERKC-C method. Noting that
@@
we immediately obtain $V(t_{k+1})=V_{k+1}$ and $V(t_{k,i})=V_{k,i}$ for $k=0,1,\ldots,N-1$ and $i=1,2,\ldots,s$. Since $\ell_{i}$ is a polynomial of degree $s-1$, the coefficients $b_{i}(\theta;-tA)$ are actually linear combinations of the functions
$\varphi_{j}(-tA)=\frac{1}{t^{j}}\int_{0}^{t}\mathrm{e}^{-(t-\sigma)A}\frac{\sigma^{j-1}}{(j-1)!}~{}\mathrm{d}\sigma,\quad 1\leq j\leq s.$ (3.14)
Recalling the parabolic smoothing property of $A$, we have the estimate
$\|t^{\gamma}A^{\gamma}b_{i}(\theta;-tA)\|_{X\leftarrow X}\leq C,\quad 0\leq\gamma\leq 1,~{}t\geq 0.$
@@
where $\Pi_{h}U(t)=(1-\theta)U_{k}+\theta U_{k+1}$, for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $10\leq k\leq n-$ 1. The ERKC-C variant is
$V_{n+1}=\mathrm{e}^{-h_{n+1}A}V_{n}+h_{n+1}\varphi_{1}(-h_{n+1}A)g(t_{n+1},V_{n+1},V^{\tau}(t_{n+1})),\$
where $V(t)=\mathrm{e}^{-\theta h_{k+1}A}V_{k}+\theta h_{k+1}\varphi_{1}(-\theta h_{n+1}A)G^{\rm C}_{k,1}$, for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $10\leq k\leq n-$ 1.

### v3

@@ Abstract
In this paper, exponential Runge–Kutta methods of collocation type (ERKC) which were originally proposed in (Appl Numer Math 53:323–339, 2005) are extended to semilinear parabolic problems with time-dependent delay. Two classes of the ERKC methods are constructed and their convergence properties are analyzed. It is shown that methods with $s$ arbitrary nonconfluent collocation parameters achieve convergence of order $s$. Provided that the collocation parameters fulfill some additional conditions and the solutions of the problems exhibit sufficient temporal and spatial smoothness, we derive superconvergence results. Finally, some numerical experiments are presented to illustrate our theoretical results.
@@ 1 Introduction
$\left\{\begin{aligned} &u^{\prime}(t)+Au(t)=g(t,u(t),u^{\tau}(t)),&&t>0,\\ &u(t)=\phi(t),&&t\leq 0,\end{aligned}\right$. (1.1)
@@
In the past, significant numerical research has been conducted for problems of the form (1.1) with constant delay; see, for example, [17, 24, 28]. Compared to constant delay models, time-dependent delay models generally provide a more accurate description of the dynamic nature of real-world systems. Nevertheless, the presence of time-dependent delays complicates both the mathematical modeling and the numerical solution, and relatively little numerical analysis has been performed for these models. Despite these difficulties, some studies have been conducted in this area. For example, Tang and Zhang [23, 32] constructed fully discrete $\theta$-methods to solve semilinear reaction-diffusion equations and Sobolev equations with time-dependent delay. However, their approach requires the absence of discontinuities in the true solution, which is a significant limitation, as discontinuities commonly arise in delay differential equations. Operator splitting for abstract delay equations has also been investigated in [1, 2, 10]. Until now, there have been limited numerical studies on temporal high-order schemes for solving (1.1) that can account for discontinuities due to the presence of a time-dependent delay.
@@
In recent years, exponential integrators have attracted considerable attention due to their effectiveness on stiff semilinear systems, such as (1.1) with $A$ having a large norm or being an unbounded operator, and the nonlinearity $g$ satisfying a local Lipschitz condition with a relatively small Lipschitz constant in a strip along the exact solution. The stiff convergence of exponential integrators, such as exponential Runge–Kutta methods [12, 13, 19], exponential Rosenbrock methods [16] and exponential multistep methods [15], has been investigated. Exponential integrators have also been extended to various areas, such as low-regularity integrators [21], exponential splitting methods [9], and phase field models [18]. For a detailed overview of exponential integrators, we refer the reader to [14]. Some research on exponential integrators for (parabolic) differential equations with constant delay can be found in [4, 6, 29, 30, 31, 33, 34]. In this paper, we aim to extend the exponential Runge–Kutta methods of collocation type (ERKC) as presented in [13] to the larger class of problems (1.1). For this purpose, we construct continuous extensions of the discrete solutions obtained by the ERKC methods, and we derive the convergence results. In contrast to the aforementioned works on exponential integrators for problems with constant delay, our research focuses on the following three specific topics.
  • (a)
    The existing works are restricted to nonlinearities of the form $g:[0,T]\times X\times X\to X$, where $X$ denotes the state space of problem (1.1), our analysis allows for a broader class of nonlinearities (see Assumption 2.2).
  • (b)
    In the constant-delay case, by choosing a uniform time step that divides $\tau$, it follows that for every mesh point $t_{n}$, the delayed argument $t_{n}-\tau$ also coincides with a mesh point. This alignment simplifies the approximation of the delayed solution $u(t_{n}-\tau)$ during time stepping. In contrast, for time-dependent delay, the delayed argument $t_{n}-\tau(t_{n})$ generally does not coincide with any mesh point. To address this, we propose two strategies for constructing continuous extensions of the numerical solution to approximate the delayed solution.
  • (c)
    The existing studies are limited to uniform time stepping. However, for problems with time-dependent delay, variable step sizes are required to ensure that the discontinuity points of the solution are captured by the time mesh. The corresponding convergence analysis on nonuniform meshes is nontrivial and requires new techniques.
@@
The outline of the paper is as follows. In Section 2, we summarize the employed abstract framework. Further, we construct two classes of $s$-stage ERKC methods for (1.1) in Section 3. The error analysis for these methods is carried out in Section 4, and an error estimate of order $s$ is obtained. Provided that the collocation parameters satisfy some additional conditions and the problem possesses sufficient temporal and spatial regularity, we obtain superconvergence results of order $s+1$ and $s+1+\beta$ (see (5.17) for the definition of $\beta$) in Section 5. Finally, we present some numerical experiments in Section 6 to illustrate the theoretical results.
@@ 2 Analytical Framework
Let $A:D(A)\rightarrow X$ be sectorial and $\overline{D(A)}=X$. Without restriction of generality, we assume that the spectrum $\sigma(A)$ of $A$, satisfies $\mathrm{Re}\,\sigma(A)>0$.
Under this assumption, the operator A-A is the infinitesimal generator of an analytic semigroup $\{\mathrm{e}^{-tA}\}_{t\geq 0}$, and the fractional powers of $A$ are well defined. We recall that $A$ satisfies the parabolic smoothing property (see [11])
$\|\mathrm{e}^{-tA}\|_{X\leftarrow X}+\big\|t^{\gamma}A^{\gamma}\mathrm{e}^{-tA}\big\|_{X\leftarrow X}\leq C,\quad\gamma,t\geq 0,$
where for two Banach spaces $X$ and $Y$, the symbol $\|\cdot\|_{Y\leftarrow X}$ denotes the operator norm from $X$ to $Y$.
For $0\leq\alpha<1$, we consider the Banach space $V=\{v\in X:A^{\alpha}v\in X\}$ with norm $\|v\|_{V}=\|A^{\alpha}v\|$. Our basic assumption on $g$ is the following.
@@
$\left\{\begin{aligned} &\frac{\partial u}{\partial t}(t,{x})-\Delta u(t,{x})=\frac{1}{1+u(t,{x})^{2}}+\frac{1}{1+u(\frac{t}{2}-\frac{1}{2},{x})^{2}},&&t\in(0,T],~{x}\in\Omega,\\ &u(t,{x})=0,&&t\in(0,T],~{x}\in\partial\Omega,\\ &u(t,{x})=\phi(t,{x}),&&t\in[-\tfrac{1}{2},0],~{x}\in\Omega,\end{aligned}\right$. (2.1)
@@
    $X=C_{0}(\Omega)$, $D(A)=\{u:u\in W^{2,p}(\Omega)~\mbox{for all}~p\geq 1,\,\Delta u\in C_{0}(\Omega),\,u|_{\partial\Omega}=0\}$.
@@ 3 ERKC methods
An important feature of delay (partial) differential equations is that they possess solutions with discontinuities in the derivative at initial time and discontinuities in high order derivatives at subsequent points. This phenomenon has been well studied in the context of delay differential equations [3]. We provide a brief explanation of how this phenomenon also arises in delay parabolic equations. Suppose that $\phi\in C^{2}([-\tau_{0},0];V)$ and $\tau\in C^{2}([0,T])$, and assume that the function $g$ is twice continuously Fréchet differentiable and that $\tfrac{\partial^{2}g}{\partial{v^{2}}}(t,v,w)$ is locally Lipschitz continuous. Within the interval $0,\xi_{1}]$, where $\xi_{1}-\tau(\xi_{1})=0$, the initial value problem (1.1) reduces to a problem without delay
$u^{\prime}(t)+Au(t)=h(t,u(t)),\quad t\in[0,\xi_{1}],\quad u(0)=\phi(0),$
where $h(t,v)=g(t,v,\phi(t-\tau(t)))$. If $u(0)\in D(A)$ and the compatibility conditions
$r_{0}=-Au(0)+h(0,u(0))\in D(A),\quad-Ar_{0}+\frac{\partial h}{\partial t}(0,u(0))+\frac{\partial h}{\partial v}(0,u(0))r_{0}\in V$
hold, we deduce from [20, Theorem 7.1.2] that the solution satisfies $u\in C^{2}([0,\xi_{1}];V)$. Noting that the compatibility conditions also hold at $t=\xi_{1}$, it follows that $u\in C^{2}([\xi_{1},\xi_{2}];V)$ as well, where $\xi_{2}-\tau(\xi_{2})=\xi_{1}$. If $\phi^{\prime}(0^{-})\neq u^{\prime}(0^{+})$ in $V$, noting that $(t-\tau(t))^{\prime}\neq 0$ due to the strict monotonicity of $t-\tau(t)$, we obtain
$u^{\prime\prime}(\xi_{1}^{-})-u^{\prime\prime}(\xi_{1}^{+})=\frac{\partial g}{\partial w}(\xi_{1},u(\xi_{1}),\phi(\xi_{1}-\tau(\xi_{1})))\big(\phi^{\prime}(0^{-})-u^{\prime}(0^{+})\big)\cdot(t-\tau(t))^{\prime}|_{t=\xi_{1}}\neq 0,$
where the notation $v(t^{\pm})$ denotes the one-sided limits of $v$ as $t\to t^{\pm}$. The same argument can be extended to show that $u:[\xi_{1},T]\to V$ is twice continuously Fréchet differentiable. The above analysis indicates that discontinuities also arise in delay parabolic equations.
The set of potential discontinuity points $J=\{\xi_{\mu}:\mu\in\mathbb{N},~\xi_{\mu}<T\}$ (which are called primary discontinuity points in the theory of delay differential equations [3]) are generated by the recursion
@@
For $\xi_{m}<T\leq\xi_{m+1}$ with some $m\in\mathbb{N}$, we have $J=\{\xi_{\mu}\}_{\mu=0}^{m}$ and corresponding intervals $\mathcal{I}^{\mu}=[\xi_{\mu-1},\min\{\xi_{\mu},T\}]$ for $\mu=1,2,\ldots,m+1$. At the points $\xi_{\mu}\in J$, the solutions of (1.1), regardless of how regular the given functions $\phi$, $g$ and $\tau$ are, will generally exhibit a low degree of regularity. In particular, if the functions $\phi$, $g$, and $\tau$ are $p$-times continuously Fréchet differentiable and $\frac{\partial^{p}g}{\partial v^{p}}(t,v,w)$ is locally Lipschitz continuous, then, provided that certain compatibility conditions are satisfied at $t=0$, it follows from [20, Theorem 7.1.2] that the solution $u\in C^{p}(\mathcal{I}^{\mu};V)$, $\mu=1,2,\ldots,m+1$. Considering such discontinuity points, we choose the mesh
@@
such that $J\subseteq I_{h}$. Choosing nonconfluent collocation parameters $c_{i}\in[0,1]$ for $i=1,2\ldots,s$, we set
$I_{n+1}=[t_{n},t_{n+1}],~~h_{n+1}=t_{n+1}-t_{n},~~t_{n,i}=t_{n}+c_{i}h_{n+1}~~\mbox{and}~~h=\max_{1\leq n\leq N}h_{n}.$
@@
$u(t_{n}+h_{n+1}=\mathrm{e}^{-h_{n+1}A}u(t_{n})+\int_{0}^{h_{n+1}}\mathrm{e}^{-(h_{n+1}-\sigma)A}g(t_{n}+\sigma,u(t_{n}+\sigma),u^{\tau}(t_{n}+\sigma))~\mathrm{d}\sigma.$ (3.1)
@@
$G_{n}(\sigma)=\sum_{j=1}^{s}\ell_{j}(\sigma/h_{n+1})G_{n,j},\quad\mbox{with}~G_{n,j}=g(t_{n,j},U_{n,j},u^{\tau}(t_{n,j})),$
where $U_{n,j}$ is an approximation to $u(t_{n,j})$ which will be determined later, and $\ell_{j}$ is the Lagrange interpolation polynomial
$\ell_{j}(\rho)=\prod_{m=1,m\neq j}^{s}\frac{\rho-c_{m}}{c_{j}-c_{m}}.$
@@
$b_{i}(-h_{n+1}A)$ $=\int_{0}^{1}\mathrm{e}^{-(1-\xi)h_{n+1}A}\ell_{i}(\xi)~\mathrm{d}\xi$ (3.3)
$=\frac{1}{h_{n+1}}\int_{0}^{h_{n+1}}\mathrm{e}^{-(h_{n+1}-\sigma)A}\ell_{i}(\sigma/h_{n+1})~\mathrm{d}\sigma.$
@@
$a_{ij}(-h_{n+1}A)$ $=\int_{0}^{c_{i}}\mathrm{e}^{-(c_{i}-\xi)h_{n+1}A}\ell_{j}(\xi)~\mathrm{d}\xi$ (3.5)
$=\frac{1}{h_{n+1}}\int_{0}^{c_{i}h_{n+1}}\mathrm{e}^{-(c_{i}h_{n+1}-\sigma)A}\ell_{j}(\sigma/h_{n+1})~\mathrm{d}\sigma.$
The scheme (3.2)-(3.5) is not computable yet, since the delayed values $u^{\tau}(t_{n,i})$ in $G_{n}$ still have to be approximated. One feasible approach is to construct a continuous approximation by using piecewise interpolation polynomials based on the known values $U_{k}$ and $U_{k,i}$ for $k=0,1,\ldots,n-1$ and $i=1,2,\ldots,s$. For this, we introduce an interpolation $\Pi_{h}U(t)$, which is a polynomial of degree $l$ when restricted to each interval $I_{k+1}$. The polynomial $\Pi_{h}U|_{I_{k+1}}$ is based on the nodes of the set $S_{k+1}=\{t_{k},t_{k+1}\}\cup\{t_{k,1},t_{k,2},\ldots,t_{k,s}\}$, satisfying
$\Pi_{h}U(t_{k})=U_{k},~\Pi_{h}U(t_{k+1})=U_{k+1},~\mbox{and}~\Pi_{h}U(t_{k,i})=U_{k,i}~~\mbox{for}~1\leq i\leq s.$ (3.6)
[…]
@@
$U_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}U_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm I}_{n,j},\quad 1\leq i\leq s,$
$G^{\rm I}_{n,j}$ $=g(t_{n,j},U_{n,j},(\Pi_{h}U)^{\tau}(t_{n,j})),\quad 1\leq j\leq s,$
@@
The scheme (3.7) is one option among several possible schemes for extending ERKC methods to (1.1). We shall introduce an alternative approach for constructing an approximation to the delayed solution $u^{\tau}(t_{n,i})$ in $G_{n}$. To distinguish it from the previous introduced numerical solution $U_{n+1}$ and $U_{n,i}$ in (3.7), we denote the approximate solutions of the new method by $W_{n+1}$ and $W_{n,i}$:
$W_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}W_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(-h_{n+1}A)g(t_{n,i},W_{n,i},W^{\tau}(t_{n,i})),$
$W_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}W_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)g(t_{n,j},W_{n,j},W^{\tau}(t_{n,j})),~~1\leq i\leq s,$
with $W(t)$ still to be defined. Instead of using piecewise interpolation polynomials, we construct a continuous solution $W(t)$ by setting $W(t)=\phi(t)$ for $t\in[-\tau_{0},0]$ and, for $t=t_{k}+\theta h_{k+1}$ with some $k\in\{1,2,\ldots,n\}$ and $\theta\in(0,1]$,
$W(t_{k}+\theta h_{k+1})=\mathrm{e}^{-\theta h_{k+1}A}W_{k}+h_{k+1}\sum_{i=1}^{s}b_{i}(\theta;-h_{k+1}A)g(t_{k,i},W_{k,i},W^{\tau}(t_{k,i})),$ (3.8)
@@
$b_{i}(\theta;-h_{k+1}A)$ $=\int_{0}^{\theta}\mathrm{e}^{-(\theta-\xi)h_{k+1}A}\ell_{i}(\xi)~\mathrm{d}\xi$ (3.9)
$=\frac{1}{h_{k+1}}\int_{0}^{\theta h_{k+1}}\mathrm{e}^{-(\theta h_{k+1}-\sigma)A}\ell_{i}(\sigma/h_{k+1})~\mathrm{d}\sigma.$
@@
$W_{n+1}$ $=\mathrm{e}^{-h_{n+1}A}W_{n}+h_{n+1}\sum_{i=1}^{s}b_{i}(-h_{n+1}A)G^{\rm C}_{n,i},$ (3.10)
$W_{n,i}$ $=\mathrm{e}^{-c_{i}h_{n+1}A}W_{n}+h_{n+1}\sum_{j=1}^{s}a_{ij}(-h_{n+1}A)G^{\rm C}_{n,j},\quad 1\leq i\leq s,$
$G^{\rm C}_{n,j}$ $=g(t_{n,j},W_{n,j},W^{\tau}(t_{n,j})),\quad 1\leq j\leq s.$
For $A=0$, this scheme reduces to a collocation method, in which the delayed solution is approximated by the collocation polynomials. For this reason, we refer to it as the ERKC-C method. Noting that
@@
we immediately obtain $W(t_{k+1})=W_{k+1}$ and $W(t_{k,i})=W_{k,i}$ for $k=0,1,\ldots,N-1$ and $i=1,2,\ldots,s$. For each fixed $\theta\in[0,1]$, the coefficients $b_{i}(\theta;-h_{n+1}A)$, as defined in (3.9), are linear combinations of the functions $\varphi_{j}(-\theta h_{n+1}A)$, where
$\varphi_{j}(-tA)=\int_{0}^{1}\mathrm{e}^{-(1-\xi)tA}\frac{\xi^{j-1}}{(j-1)!}~\mathrm{d}\xi=\frac{1}{t^{j}}\int_{0}^{t}\mathrm{e}^{-(t-\sigma)A}\frac{\sigma^{j-1}}{(j-1)!}~\mathrm{d}\sigma,\quad 1\leq j\leq s.$ (3.12)
The coefficients in the linear combination depend on $\theta$; see, e.g., Example 3.2. Recalling the parabolic smoothing property of $A$, we have the estimate
$\|t^{\gamma}A^{\gamma}b_{i}(\theta;-tA)\|_{X\leftarrow X}\leq C,\quad 0\leq\gamma\leq 1,~t\geq 0.$ (3.13)
@@
where $\Pi_{h}U(t)=(1-\theta)U_{k}+\theta U_{k+1}$, for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $0\leq k\leq n-1$. The ERKC-C variant is
$W_{n+1}=\mathrm{e}^{-h_{n+1}A}W_{n}+h_{n+1}\varphi_{1}(-h_{n+1}A)g(t_{n+1},W_{n+1},W^{\tau}(t_{n+1})),\\$
where $W(t)=\mathrm{e}^{-\theta h_{k+1}A}W_{k}+\theta h_{k+1}\varphi_{1}(-\theta h_{n+1}A)G^{\rm C}_{k,1}$, for $t=t_{k}+\theta h_{k+1}\in I_{k+1}$, $0\leq k\leq n-1$.